OpenStudy (anonymous):

Solve and describe the steps you took to solve: |m-2|-6=-2

OpenStudy (mathstudent55):

To solve an absolute value equation like yours, you need the absolute value part isolated. The first step is to add 6 to both sides. Do that first. What do you get?

OpenStudy (anonymous):

4?

OpenStudy (mathstudent55):

You are correct. The right side is 4, but you need to write the entire equation: |m - 2| - 6 = -2 Add 6 to both sides to get: |m - 2| = 4 Ok?

OpenStudy (anonymous):

Ok.

OpenStudy (mathstudent55):

Now we do the step to take care of the absolute value part. When you have an expression inside absolute value signs equaling a number, to get rid of the absolute value signs, you turn the equation into two equations separated by the word "or". The two equations are (1) the expression inside the absolute value signs equaling the number, and (2) the expression inside the absolute value signs equaling the negative of the number. The expression you have is simply m - 2, so you get this: m - 2 = 4 or m - 2 = -4 Now you solve each simple equation and separate the two answers with the word or. That is the solution to the eqaution.

OpenStudy (anonymous):

Ah, I see. Okay, thank you.

OpenStudy (mathstudent55):

Here's the whole solution: |m - 2| - 6 = -2 Add 6 to both sides to get: |m - 2| = 4 Separate into two simple equations separated by "or". m - 2 = 4 or m - 2 = -4 Add 2 to both sides in both equations: m = 6 or m = -2 This is the final solution.